OpenStudy (anonymous):

The following differential equation describes a periodically forced, harmonic (undamped) oscillator: x¨ + ω0^2 x = A cos ωt, where ω0, A and ω are positive real constants. i) Calculate the homogeneous and particular solutions for this equation and hence write down the general solution. ii)) Describe what happens to x as ω → ω0. What is this phenomenon called

OpenStudy (unklerhaukus):

ii) resonance

OpenStudy (unklerhaukus):

\[\ddot x + \omega_0^2 x = A \cos \omega t\]

OpenStudy (unklerhaukus):

the general solution is the sum of the particular solution and the corresponding homogenous equation \[x_c(t)+x_p(t)=x(t)\] the corresponding homogenous equation is \[\ddot x_c + \omega_0^2 x_c =0\]

OpenStudy (unklerhaukus):

the characteristic auxiliary equation is \[m^2+\omega_0^2=0\] \[b^2-4ac=0^2-4\times 1\times \omega_0^2=-(2\omega_0)^2<0\]

OpenStudy (unklerhaukus):

\[\frac{-b\pm\sqrt{-(2\omega_0)^2}}{2a}=\frac{\pm 2\omega_0\sqrt{-1}}{2}=\pm i\omega_0\]

OpenStudy (unklerhaukus):

@nadian63

OpenStudy (anonymous):

Thank you!

OpenStudy (unklerhaukus):

what did you get as your your final solution for \(x(t)\)

OpenStudy (anonymous):

I need to finish it